OpenStudy (16silly):

Solve the quadratic equation m2+ 2m - 24 = 0 by factoring.

OpenStudy (mathstudent55):

Do you know how to factor?

OpenStudy (16silly):

I think I was taught one but I don't remember.

OpenStudy (dogzcatz):

m2+ 2m - 24 = 0 4m-24=0 4m=24 m=24÷4 This is the only way I know how to solve it sorry

OpenStudy (16silly):

I think that the question messed up. The first m2 is supposed to be m^2.

jhonyy9 (jhonyy9):

m^2 +2m -24 =0 solve by factoring mean that you need getting two numbers with sum equal 2 and product equal -24 - can you get them ?

OpenStudy (16silly):

No.

OpenStudy (mathstudent55):

To factor a quadratic trinomial of the type x^2 + ax + b = 0, find two numbers whose product is b and whose sum is b.

jhonyy9 (jhonyy9):

you know that x+y=2 x*y=-24 can you calculi the x and y ?

OpenStudy (mathstudent55):

Let's look at your problem. To factor m^2+ 2m - 24 = 0 find two numbers that multiply to -24 and add to 2.

OpenStudy (mathstudent55):

Can you list several sets of two numbers that multiply to -24? Here is an example: -24, 1 Can you list other sets of two numbers that multiply to -24?

OpenStudy (mathmale):

Yes. In other words, look at the constant term, -24, and determine its factors. Which of these two factors combine to give you the middle term, 2m? Please show your work. Factors of -24 are ... ?

OpenStudy (16silly):

2,12, 4,6

OpenStudy (16silly):

None of those add up to 2

OpenStudy (16silly):

nvm I figured it out. Its -4, 6

jhonyy9 (jhonyy9):

@16silly do you know negativ numbers ?

OpenStudy (mathstudent55):

Good. Since the product has to be negative, each set of numbers yields two results: Here are all the possible sets of two integers that multiply to -24 -24, 1 -1, 24 -2, 12 -12, 2 -4, 6 -6, 4 -3, 8 -8, 3

OpenStudy (16silly):

I got it. it is -4 and 6

OpenStudy (mathmale):

Now factor your equation. Check your work by substituting -4 and 6. Is the equation true when you do this (one value at a time)?

OpenStudy (mathstudent55):

Now you add both numbers in each set, and choose the one that adds to 2. You already did it. -4, 6 Good.

OpenStudy (16silly):

I forgot it was -24

OpenStudy (mathstudent55):

To factor, you set up two sets of parentheses, and fill in the variabkle, m, in each: (m )(m ) = 0 What goes int he blanks is the numbers you found, so you get: (m + 6)(m - 4) = 0 The order does not matter because of the commutative property of multiplication. You can also write (m - 4)(m + 6) = 0

OpenStudy (mathstudent55):

The factoring part is done. Do you understand it so far?

OpenStudy (16silly):

Yes

OpenStudy (mathstudent55):

Since you have a product that equals zero, you set each factor equal to zero and solve each equation. (m - 4)(m + 6) = 0 m - 4 = 0 or m + 6 = 0 Solve each equation for m. The solution is both of the values you get separated by the word "or".

OpenStudy (mathmale):

@16silly: what are the solutions? List them, please.